OpenStudy (anonymous):

Identify the type of conic section that has the equation 9x^2+ 16y^2 = 144 and identify its domain and range.

OpenStudy (anonymous):

A parabola is in the form\[ 4py=x^2 \] An ellipse is in the form\[ \left(\frac{x}{a}\right)^2+\left(\frac{y}{b}\right)^2=1 \] A hyperbola is in the form\[ \left(\frac{x}{a}\right)^2-\left(\frac{y}{b}\right)^2=1 \]

OpenStudy (anonymous):

oh.. so i would solve for those values and plug them in?

OpenStudy (anonymous):

Nope. Just think of which one your equation looks most similar to

OpenStudy (anonymous):

oh.. well i would say the second one. is that right?

OpenStudy (anonymous):

Yeah, so what is it called?

OpenStudy (anonymous):

a hyperbola?

OpenStudy (anonymous):

... no it says ellipse.

OpenStudy (anonymous):

the second one does?

OpenStudy (anonymous):

"An ellipse is in the form" This is me saying "this second equation here is for an ellipse"

OpenStudy (anonymous):

oh.. got it. sorry, I found this really confusing. So then the domain would be all real numbers right?

OpenStudy (anonymous):

No, the domain is between \([-a, a]\) and the range is between \([-b,b]\)

OpenStudy (anonymous):